plese graph the hyperbola.
(x^2)/(25)-(y^2)/(144)=1

Answer by stanbon(57967)

plese graph the hyperbola.
(x^2)/(25)-(y^2)/(144)=1
------------------------------
Since that equation is not a function, this site
cannot show a graph of the hyperbola.
----
You could plot a few points then sketch the double curve.
x-intercepts:
If y = 0, x^2 = 25, so you have points (5,0) and (-5,0)
---------------------------------
Note that if x = 0, y^2 = 144 so y is not a Real Number
That means there are no y-intercepts.
-------
The hyperbola opens to the left and to the right from
those x-intercepts points.
========================================
Cheers,
Stan H.